OpenStudy (anonymous):

Need help with pre-cal review 1. Write the equation of a parabola with vertex (1, -9) and directrix y = -16. 2. What is the focus of the equation that you found in problem #1? 3. What is the axis of symmetry of the equation you found in problem #1?

OpenStudy (sweetburger):

i swear your book has the formula's to find each of these

OpenStudy (anonymous):

Vertex form of a parabola is Y=a(x-h)^2+k h and k are the vertex So we only need to solve for a Knowing the directrix is -16 and the y value of the vertex is -9, we find the difference of those to be 7, so we add 7 to -9 and that gives us the value of a. ^_^

OpenStudy (anonymous):

So would it be -9= -2(x-1)^2 + -9?

OpenStudy (anonymous):

Ya, but replace -9= with y=

OpenStudy (anonymous):

so the equation is y=-2(x-1)^2+(-9)?

OpenStudy (anonymous):

Ya, but u don't need the () around the -9

OpenStudy (anonymous):

I was just making sure the negative was obvious. The format in my text book is different though. The equations are set up like (y-k)^2=4p(x-h).

OpenStudy (anonymous):

So it'd be like (y+9)^2=4p(x+1). What is "p" though?

OpenStudy (anonymous):

For part 2, remember the -2 u got for a? That is the y value for the focus. And the x value for the focus is the x value from the vertex

OpenStudy (anonymous):

I think their p is the same as our a

OpenStudy (anonymous):

P is the distance between the focus and vertex

OpenStudy (anonymous):

That equation from the book looks like it's for parabolas that open sideways not vertical...

OpenStudy (anonymous):

Yup, they're sideways.

OpenStudy (anonymous):

Yayaya, then p is the same as our a

OpenStudy (anonymous):

So (y+9)^2=-8(x-1)?

OpenStudy (anonymous):

If your equation for Our parabola given supposed to be open sideways then our directrix should be x equals not Y equals

OpenStudy (anonymous):

Yes that must be correct

OpenStudy (anonymous):

No, the book says if it opens upward (vertical) then it'd be x equals

OpenStudy (anonymous):

Alright then. That's different then what I learned but u do what looks right

OpenStudy (anonymous):

Haha. Thanks for your help. :)

OpenStudy (anonymous):

^_^